OpenStudy (anonymous):

angle_1and angle_2 are supplementary angles. mangle_1 is 4y + 7 and mangle_2 is 9y + 4. Find mangle_2.

OpenStudy (anonymous):

answer choices?

OpenStudy (anonymous):

a supplementary angle adds up to 180. 4y+7 = 180 angle 1 y = 43.25 9y + 4 = 180 y = 19.6

OpenStudy (anonymous):

angle_1and angle_2 are supplementary angles. m angle_1 is 4y + 7 and m angle_2 is 9y + 4. Find m angle_2. @alexxxis @ninab731

OpenStudy (anonymous):

a supplementary angle is one that adds up to 180 degrees

OpenStudy (anonymous):

Supplementary angles combine to equal 180 degrees. so 13y+11=180

OpenStudy (anonymous):

Just solve for y and substitute your y-value back into the equation for the measure of angle 2

OpenStudy (anonymous):

169=13^2